11 May 2015 MODS2 Commissioning

Observers: Doug Miller, Mark Wagner, Olga Kuhn
OSA: Steve Allanson
InstSp: Jenny Power

Photometric Skies. Seeing is good, 0.6" on guider, 0.55" on WFS.

Summary of issues:
  • One through 120-sec slit image is smeared. Probably poor IMCS-lock. We saw a similar thing for MODS1 and this is discussed in an email from Rick dated 12/02/2013 in the thread "reference position for modsAlign -r". Was a problem with threshold - IMCS thought it was locked when it wasn't. But, this through-slit image was not the first after instconfig but the last (the 12th of 12 images), it may be simply that too much time/elevation had elapsed since the instconfig, and the closed-loop imcslock made some needed adjustments to the collimator. But the fainter portion of the slit image (the shadow) is aligned with the previous slit image; if this explanation were correct, one would expect the opposite.
  • imcslock times out when changing from blue-only grating to dual prism (this happened 3x).
  • rightzeropoint found to be off --- with initial value of 27.55, line of stars was tilted CCW from the slit. Adjusted it by 0.6 deg. Images r 41-51 are with initial value, images r 69-79 are with current value of 26.95.
  • unusual saturation effects on red channel (see r 83 and 85 for examples).
  • stray light - as expected for adsec.
  • series of ~6-7 blue images with sequencer problems (42,44-47 at least). Restart of Blue IC done, after which the next two blue images were OK, but the 3rd (51) is again bad.

Doug took some on-axis WFS images. Guider HS and pupil position both look good --- the same as last time we operated MODS2 AGw.

03:00:00:
  • acqMODS2 Hz43.
  • mods2Align mods2r.20150511.0002.fits
    • MODS Offset Command:
         offsetxy -2.261  6.774 rel
      Send the offset to the Telescope (Y|N)?: y
         Sending offset...
      *** ERROR: Command 'offsetxy -2.261 6.774 rel' timed out after 60 seconds
         Offset command returned error - /home/MODSeng/bin/mods2Align aborting
      
      
  • sent offset from MODS2 command window. This worked fine. Unsure why the error. Looks well centered after the offset.
  • Hz43 has a companion 3" away --- the PA between the two stars is about 90 deg, yet the script uses a PA = 45 deg and both stars are in the slit, yet with a reduced Y-separation. The alignment looks good, though I may have chosen the wrong star. Due to c confusion and small Y-separation between spectra, moving to a different standard which is close by, GD153.

03:15: Doug noticed that pupil shift was not what was expected for the off-axis guide star. He's looking into it. (It was only that his IDL program was not using most up-to-date values that were in GCS. Once changed, they are OK.)

03:20 acqMODS2 ~modseng/modsScripts/modsSpecPhot/gd153.acq
  • timeout again. offset command does not appear to go through the GUI: dX, dY not changing and Send Offset button not orange.
  • messages:
MODS Offset Command:
   offsetxy -2.231  7.921 rel

Send the offset to the Telescope (Y|N)?: Y
   Sending offset...
*** ERROR: Command 'offsetxy -2.231 7.921 rel' timed out after 60 seconds
   Offset command returned error - /home/MODSeng/bin/mods2Align aborting
  • skyping Rick. He made a change.

03:45 in meantime, taking a slitscan. We see some misalignment between slit and detXY Y-axis.

04:00 acqMODS2 gd153.acq. OK now. execMODS2 to observe star in all 6 spectroscopic modes.
  • dual grating: b 2-4 and r 14-16
  • red-only grating: r 17-19
  • blue-only grating b 5-7
  • dual-prism --- IMCS timed out, quickly locked upon "R"etry. b

05:15 acqMODS2 hz44.acq
  • offsetxy -1.906 3.430 rel

05:15 execMODS2 hz44.obs to observe at all 6 modes. AM=1.002

05:40 did a quick calculation of angle of detxy y-axis wrt slit. Is right rotator zero point OK?
  • detxy y-axis is -0.676 deg from +Y (CCW)
  • slit axis is -0.132 deg from +Y
  • detxy y-axis to slit is -0.544 ... we want to rotate slit CCW by 0.544 deg, and stars CW by 0.544. Subtract 0.544 from right rotator zeropoint (27.55). Should be 27.00. Will rerun script, redo calc based on those data and adjust zeropoint and confirm later.
  • again, going from blue-only grating to dual prism, blue imcslock took longer than the timeout.

06:28 execMODS2 GD153act.img to collimate on GD153, prior to running the script to take images at a range of offsets along Y: PRESET TRACK
offsetxy 0 0 abs
offsetxy 0 30 abs
offsetxy 0 -10 abs
offsetxy 0 10 abs
offsetxy 0 -30 abs
offsetxy 0 60 abs
offsetxy 0 -60 abs
offsetxy 0 20 abs
offsetxy 0 -20 abs
offsetxy 0 40 abs
offsetxy 0 -40 abs

06:30: TCS error. Steve had to reset shell. This was similar to the errors they had last night. (IT 5559)

06:42: execMODS2 GD153.img
>> offsetxy 0 30 abs
** ERROR: offsetxy IIF error - OffsetPointing result status: Error, ERROR: No result available yet
** Abort, Retry, or Ignore? > r
** Retrying command 'offsetxy 0 30 abs'

Stars (detxy Y-axis) tilted wrt detector by -0.788 deg (CCW). Slit measurement (0052) is affected by poor imcslock(?) (seen this before for mask alignments where all boxes are "shadowed"). Compare 0052 with 0011 (0.6" slit taken earlier tonight) and the smearing is readily visible.

Use old slit (0011), which is tilted wrt detector by -0.132 deg.

Then line of stars needs to be rotated CW by 0.65 deg.

Combined with earlier results, let's take 0.6 deg and adjust rotator zero point by -0.6 deg (27.55 - 0.6 = 26.95 deg).

07:02: acqMODS2 gd153.acq.
Log file _modsalign.log open
#   COL    LINE   COORDINATES
#     R    MAG    FLUX     SKY    PEAK    E   PA BETA ENCLOSED   MOFFAT DIRECT
 524.57  525.68 524.57 525.68
  13.80  10.09 920190.    357.  29223. 0.03  -66 3.93     4.58     4.63   4.60

Computed Slit Alignment Offset:
   dX:  -2.456 arcsec
   dY:   7.911 arcsec


MODS2 Offset Command:
   offsetxy -2.456  7.911 rel
Through-slit confirmatory image (54) is not well centered this time. Requires 0.8" dx.

07:15 execMODS2 gd153.obs airmass = 1.17-1.46.
  • comm drop on blue: blue expdone needed at end of script (EXPSTATUS DON)
  • blue fitsflush done

08:32 edited RIGHTZEROPOINT from 27.55 to 26.95 in PCSInstrument.conf file. Steve restarts PCS. To confirm:
  • execMODS2 GD153act.img, followed by
  • execMODS2 GD153.img
  • note --- GD153 always seems to appear in a different place on initial active preset. Is this because of the updatepointing both the active and track preset scripts?
  • images mods2r.20150511.N.fits N=0069 - 0079. The line of stars is at an angle -0.229 deg from the Y axis. The slit was measured to be at an angle -0.13 -- -0.16 deg. Better, although not perfect. A better analysis can be done still.
  • before: r 41-51 (initial rightzeropoint of 27.55) and after: r 69-79 (current rightzeropoint 26.95)
  • see attachment (rightrotatorzp.jpg). Green points & blue line are with initial value. Yellow points and cyan line are with 26.95 deg. (ignore horizontal lines).

09:10 Astrometry on Stone L (subfield for mask2). ugriz. Seeing 0.9" on guider. Airmass 1.22. * execMODS2 SciAstrom/mods.7.StoneLf2.img
  • unusual saturation effects on red channel. See image 83.

09:20 Measurement of Rotator Center
  • execMODS2 RotCen/StoneLf2_act0.img
  • execMODS2 StoneLf2_trk180.img
    • blue image has sequencer problem (b 42).
    • red image shows saturation effect (this time on UR quad) and stray light.
  • execMODS2 StoneLf2_act180.img --- chosen guide star may have been a double star.
  • execMODS2 StoneLf2_act0.img (but I edited guide star to use one which did not cause vignetting of the MODS FOV).
    • b 43 and r 86
  • execMODS2 StoneLf2_trk180.img
    • b 44 (again a sequencer error) and r 87
  • execMODS2 StoneLf2_trk180.img again.
    • took a while to lock (repeated script and therefore instconfig command).
    • comm drop on red exptime (red exptim).
    • b45 also has a sequencer problem
  • repeat - this time just by "go" command.
    • b46 also...
  • execMODS2 StoneLf2_trk90.img (let's try another PA... at least for something new on red).
    • b47 also bad.

10:00 Steve restarted blue IC program. Will try again to collect data to measure RotCenter.
  • execMODS2 StoneL2f_act0.img
    • b 49 is OK, clear of seq problems. r 92
  • execMODS2 Stone_L2f_trk180.img
    • red imcs lock failed within timeout period, after "R"etry, achieved lock quickly.
  • execMODS2 Stone_L2f_trk90.img * b 51 --- again a sequencer problem. Repeat with "go" -b 52 is OK.
  • execMODS2 Stone_L2f_trk270.img
    • b 53 and r 96.

-- %USERSIG{OlgaKuhn - 2015-05-11}%

Comments

 
Topic revision: r4 - 11 May 2015, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback